# Cool-Core Destruction in Merging Clusters with AGN Feedback and Radiative Cooling

### v1

@@ 3.2.2 Mergers with Mass-Ratio $R=3$
Figure 4: Entropy slices on the $z=7.5$ plane for simulation $mcf5$ ($R=3$, $b=$ 464.43 kpc). The epochs are t = 0, 1, 2, 3, 4, 5, 6, 8, 10 Gyr. Each panel is 5 Mpc on a side. The cyan dot represents the SMBH of the primary cluster, while the red dot represents the SMBH of the secondary cluster (note: these dots almost overlap at 5 Gyr). The CC of the primary cluster is transformed into a NCC at the end of the simulation.
@@
The off-axis cases exhibit a more complex evolution. For the $R=3$, $b=464.43$ kpc simulation ($mcf5$), the central entropy gradually rises during the merger, peaking at $S_{40}\approx$ 470 keV cm2 between 3 and 4 Gyr. Following the SMBH merger, $S_{40}$ remains at $\approx$ 400 keV cm2 until 10 Gyr. Figure 4 shows the 2D entropy maps for $mcf5$. The first core passage results in a sideswiping motion between the two clusters, increasing the gas circular velocity and enhancing the mixing of gases with different entropies. After the SMBHs merge at 5.4 Gyr, the circular motion of gas enables further heating to maintain the central entropy at a high level. Eventually, the cluster becomes a NCC with $S_{40}\approx$ 380 keV cm2. Compared with the merger-only simulation of this case ($m5$), the $S_{40}$ evolution is roughly the same except for the peak at 3-4 Gyr. Note that after $\approx$ 8 Gyr, the central entropy starts to drop gradually due to cooling. This general feature is found in all nine $mcf$ simulations after mergers.
Figure 5: Zoom-in entropy slices at the $z=7.5$ plane for simulation $mcf5$. The epochs are t = 3.1, 3.2, 3.3, 3.4, 3.5, 3.6 Gyr. Each panel is 0.7 Mpc on a side. The high entropy near the center is due to pronounced AGN feedback.
@@
To understand the detailed evolution between 3 and 4 Gyr, we zoom in on the primary cluster core and present the entropy slices every 0.1 Gyr during 3.1-3.6 Gyr, as shown in Figure 5. During this period, the primary cluster core has not been fully disturbed by the secondary cluster. To regulate the CC, strong jet outbursts occur, raising the central entropy. Additionally, the secondary cluster falling back toward the primary cluster effectively draws gas into the primary core, providing fuel for strong AGN feedback. The effectiveness of AGN feedback is further discussed in Section 3.3.
For the $R=3$, $b=932.28$ kpc simulation ($mcf6$), the central entropy evolution is similar to that of the $mcf5$ simulation discussed above. However, the gas mixing is less effective, leading to a less significant increase in central entropy. Following the SMBH merger, $S_{40}$ gradually decreases due to cooling. Nevertheless, the final remnant is a NCC cluster, with $S_{40}\approx$ 170 keV cm2.
3.2.3 Mergers with Mass-Ratio $R=10$
@@
Figure 6: Entropy slices on the $z=7.5$ plane for simulation $mcf9$ ($R=10$, $b=$ 932.28 kpc). The epochs are t = 0, 1, 2, 3, 4, 5, 6, 8, 10 Gyr. Each panel is 6 Mpc on a side. The cyan dot represents the SMBH of the primary cluster, while the red dot represents the SMBH of the secondary cluster. The primary CC is barely disrupted during merger.
@@
3.3 Heating and Cooling
In this section, we investigate the impact of AGN feedback and radiative cooling on CC destruction in cluster merger simulations. We analyze the evolution of jet power and the accumulated energy injected by AGN feedback for each $mcf$ simulation. These parameters are then compared with the cooling rate and radiative energy to evaluate the balance between heating and cooling mechanisms.
Figure 7: Heating and cooling rate evolution for each $mcf$ simulation. Left: $R=1$ simulations. Center: $R=3$ simulations. Right: $R=10$ simulations. The solid lines are the jet power, representing the heating rate, and the dashed lines are the X-ray luminosity within 50 kpc, representing the cooling rate. The gray lines are from the single cluster simulation for comparison.
Figure 8: Evolution of accumulated heating and cooling energies for each $mcf$ simulation. Left: $R=1$ simulations. Center: $R=3$ simulations. Right: $R=10$ simulations. The solid lines are the accumulated injected AGN energy, and the dashed lines are the accumulated radiative energy. The gray lines are from the single cluster simulation for comparison.
Figure 7 illustrates the evolution of heating and cooling rates, with three subplots, each displaying three cases with the same mass ratio $R$. Simulations with the same impact parameter $b$ are plotted in the same color for comparison. Additionally, we include the result of the single cluster simulation $scf$ for reference. In each case, the solid line represents the jet power (i.e., the heating rate), and the dashed line represents the X-ray luminosity within 50 $L_{x}(<50~{}\text{kpc})$, (i.e., the cooling rate). The X-ray luminosity $L_{x}(<r)$ is calculated by
[…]
@@ 3.3 Heating and Cooling
Figure 8 shows the evolution of accumulated injected (i.e., heating) and radiative (i.e., cooling) energies. The injected energy is calculated by multiplying the jet power by each simulation time-step and summing the results. Similarly, the radiative energy is calculated by multiplying the X-ray luminosity within 50 $L_{x}(<50~{}\text{kpc})$, by each simulation time-step and summing these energies. The radius of 50 kpc for the cooling rate and radiative energy is chosen to roughly balance the levels of heating rate and injected energy.
@@
For the $scf$ simulation, the jet power increases and peaks around 1.8 Gyr to balance with the cooling rate. At this point, the accumulated injected energy approximately balances the radiative energy as well. Subsequently, the heating and cooling rates periodically reach peaks together, indicating that the system is self-regulated. In the merger simulations, the first jet power peak (determined by the initial cooling time) happens to coincide with the first core passage (determined by the initial separation and relative velocity) at $\approx 1.6-2.0$ Gyr. We examine the effect of this timing coincidence by carrying out an additional run of $mcf1$ with an initial separation between the two clusters increased by a factor of 1.5. This approach effectively delays the collision, allowing AGN feedback more time to act. We maintain a constant initial total energy, ensuring that the merging velocity between the two clusters remains consistent with the original $mcf1$ run. In this test, the first core passage occurs at 4.0 Gyr, and the SMBHs merge at 6.2 Gyr, thus delaying the evolution by 2.4 Gyr with respect to $mcf1$. During 2-4 Gyr, the jet power grows strong to balance cooling but does not significantly elevate the central entropy. The evolution of $S_{40}$ is similar to that of $mcf1$, reaching a peak after the first core passage, oscillating during the merger, and gradually decreasing post-merger. At later stages, the NCC is transitioning back to a CC. Therefore, we conclude that the results of $mcf1$ are robust and are not affected by the coincidence between cooling and the first core passage.
@@
For the $mcf1$ simulation, AGN feedback remains quiet after the first core passage until $\approx$ 7 Gyr. During this period, the radiative energy is an order of magnitude higher than the injected energy. This weak AGN feedback suggests that heating from merger is the dominant mechanism in balancing cooling. As inferred from Figure 2 and Figure 3, the merged cluster reverts to a CC state at late times, and gas mixing induced by the merger is insufficient to further heat the cluster. Consequently, the jet power rises again at the later stages to regulate cooling, and the evolution of heating and cooling becomes similar to that of the single cluster case.
@@
For the off-axis cases with $R=1$ and $R=3$ (i.e., $mcf2$, $mcf3$, $mcf5$, $mcf6$), the jet power peaks consistently during 2-4 Gyr. This occurs because in off-axis cases, the disruption from the secondary cluster is not prominent at the first core passage. Consequently, AGN feedback intensifies to balance cooling during this stage. As mentioned in Section 3.2.2, the circular motion draws gas into the central region of the primary cluster, maintaining high central density and therefore strong jet power, leading to overheating. As a result, the injected energy in these cases exceeds the radiative energy. After SMBH mergers, the clusters transform into NCCs, and both the heating and cooling rates become low.
@@
4 Discussion
4.1 Comparison with Previous Works
While there have been no other detailed cluster merger simulations including radiative cooling and AGN feedback, here we compare the results of our $mcf$ simulations with recent simulation results of Valdarnini & Sarazin (2021), which include gas cooling, star formation, and energy feedback from supernovae. Their adiabatic runs are identical to the merger simulations in ZuHone (2011), despite employing the smoothed particle hydrodynamics (SPH) scheme instead of the mesh-based scheme used in the latter. The results of their radiative simulations are generally similar to our $mcf$ simulations in terms of CC destruction. In $R=1$ off-axis cases, the final remnants are NCC clusters even if cooling is included. This is due to the continuous gas sloshing and mixing during the later stages of mergers. For the $R=10$ cases, the entropy increase is smaller compared to the merger-only runs, thus the initial CCs are preserved, especially in the off-axis cases.
The only different case is the $R=1$ head-on merger. In the simulation of Valdarnini & Sarazin (2021), this run results in a NCC cluster.  [^2] They ran another simulation with a low mass cluster (0.1$\times$ the total mass of their C1 cluster), and the CC is preserved. This suggests that the $R=1$ head-on merger case is sensitive to the initial conditions. In contrast, in our $mcf1$ simulation, the CC is heated to a NCC temporarily, then returns to a CC before the end. In addition to the difference between supernova feedback and AGN feedback (e.g., variations in the amount of energy injection, AGN jets are bipolar while supernova feedback is isotropic), there are two possible reasons for the different results. First, our CC appears to be more resilient to heating. The initial cooling time of their cluster at $0.01r_{200}$ is $\sim$2.2 Gyr, whereas the cooling time of our cluster at $0.01r_{200}$ is $\sim$1.2 Gyr. Second, there might be a significant amount of numerical heating in their simulations. They introduced a frictional term to maintain the stability of the CC entropy profile when $t<0$ and turned it off at $t=0$ to avoid affecting entropy mixing during mergers. However, this leads to numerical heating at $t>0$, resulting in an overestimation of the entropy increase. The authors suggest that this effect is more pronounced in merger simulations with large angular momentum, as the direct collision occurs later. Nevertheless, the other cases are also likely affected to some extent without the frictional term.
4.2 CC/NCC Transitions
@@
(1) CCs are preserved in minor mergers or mergers that do not trigger sufficient heating. In our $R=10$ simulations, the secondary cluster does not significantly disrupt the primary core, thereby failing to increase central entropy and destroy the CCs. In cases with insufficient heating, such as our $mcf1$ simulation, although it is heated to a NCC at the time of SMBH merger, the central cooling time is $\sim 2$ Gyr, so that the cluster eventually returns to the CC state at the end of the simulation. In this scenario, the injected AGN energy roughly balances the radiative energy (Figure 8). Therefore, the influence of mergers is subdominant, and AGN feedback plays a critical role for balancing cooling and preventing the clusters from undergoing cooling catastrophe. This is also consistent with the observational evidence that the bulk of CCs appear to be preserved over the past $\sim$ 10 Gyr (McDonald et al. 2017).
@@
(3) CCs are destroyed by the combined effects of mergers and AGN feedback. In major mergers with a large impact parameter (e.g., $mcf2$, $mcf3$, $mcf5$, $mcf6$), because direct core collisions occur at later stages, AGN feedback is required to prevent the core from catastrophic cooling. Driven by the mergers, the gas flowing into the primary core further fuels the maintenance of strong jet power, leading to overheating and CC destruction. As a result, the injected energy in these cases exceeds the radiative energy. This scenario represents an interesting case where mergers and AGN feedback work hand-in-hand to balance cooling and destroy the CCs.
@@ 4.3 Limitations of the Simulations
5 Conclusions
In this work, we conduct binary cluster merger simulations incorporating AGN feedback and radiative cooling to investigate the destruction of CCs. Our primary objective is to assess the impact of cluster mergers and AGN feedback on heating the cluster cores. We implement a purely kinetic jet feedback model mainly coupled with chaotic cold accretion, which has been shown to solve the cooling flow problem (Gaspari et al. 2013), together with an exact cooling scheme into the GAMER code. We conduct three sets of simulations: the $m[1-9]$ simulations are mergers only; the $mc[1-9]$ simulations include mergers with cooling but no AGN feedback; the $mcf[1-9]$ simulations involve mergers with both cooling and AGN feedback. Additionally, we vary the mass ratio and impact parameter to examine the entropy evolution of different merger scenarios. The important findings of our study are summarized below.

### v3

@@ III.2.2 Mergers with Mass-Ratio $R=3$
Figure 4: Entropy slices on the z=0z={0} plane for simulation $mcf6$ ($R=3$, $b=$ 932.28 kpc). The epochs are t = 0, 1, 2, 3, 4, 5, 6, 8, 10 Gyr. Each panel is 5 Mpc on a side. The cyan dot represents the SMBH of the primary cluster, while the red dot represents the SMBH of the secondary cluster (note: these dots almost overlap at 5 Gyr). The $C$ of the primary cluster is transformed into a NCC at the end of the simulation.
The off-axis cases exhibit a more complex evolution, and the results are sensitive to the interplay among merger, cooling and heating. For the $R=3$, $b=464.43$ kpc simulation ($mcf5$), the core of the primary cluster remains at the CC state ($S_{40}\lesssim$ 50 keV cm2) throughout the simulation. Unlike its non-radiative counterpart ($m5$), the merger is unable to counteract radiative cooling and transform the CC into a NCC. However, the mixing provided by merger in this case is still enough to prevent the formation of large amount of cold gas near the center. Therefore, the activity of the central AGN remains low, though still required to maintain the CC and prevent it from the cooling catastrophe.
@@
For the $R=3$, $b=932.28$ kpc simulation ($mcf6$), the central entropy gradually rises during the merger, reaching $S_{40}\approx$ 300 keV cm2 by the time of SMBH merger. Following the merger, $S_{40}$ slightly decreases until 10 Gyr due to radiative cooling. Figure 4 shows the 2D entropy maps for $mcf6$. The first core passage results in a sideswiping motion between the two clusters, increasing the gas circular velocity and enhancing the mixing of gases with different entropies. The overall evolution is similar to $mcf5$; however, the merger-induced heating is less effective due to the larger initial impact parameter. During t=34t=3-4 Gyr, there is a surge of cold gas stream feeding the central SMBH (consistent with a chaotic cold accretion-like episode), triggering jet activities that elevate the core entropy. After the SMBHs merge at 5.4 Gyr, the circular motion of gas enables further heating to maintain the central entropy at a high level. Eventually, the cluster becomes a NCC with $S_{40}\approx$ 300 keV cm2. Compared with the merger-only simulation of this case ($m6$) in Figure 2, the $S_{40}$ evolution is roughly the same, suggesting that the AGN heating has contributed to balance cooling.
Figure 5: Zoom-in entropy slices at the z=0z={0} plane for simulation $mcf6$. The epochs are t = 3.1, 3.2, 3.3, 3.4, 3.5, 3.6 Gyr. Each panel is 0.7 Mpc on a side. During this period, the central SMBH is fed by a stream of cold gas, which triggers elevated AGN activities. As a result, the core is preheated by the AGN before subsequent mixing brought by the fallback of the secondary cluster at later times.
@@
To understand the detailed evolution between 3 and 4 Gyr, we zoom in on the primary cluster core and present the entropy slices every 0.1 Gyr during 3.1-3.6 Gyr, as shown in Figure 5. During this period, the primary cluster core has not been fully disturbed by the secondary cluster. The infalling cold gas stream feeds the central SMBH and triggers the AGN activity, raising the central entropy. In other words, the AGN outbursts preheat the CC and prepare it for subsequent mixing during the fallback of the secondary cluster. The effectiveness of AGN feedback is further discussed in Section III.3.
III.2.3 Mergers with Mass-Ratio $R=10$
@@
Figure 6: Entropy slices on the z=0z={0} plane for simulation $mcf9$ ($R=10$, $b=$ 932.28 kpc). The epochs are t = 0, 1, 2, 3, 4, 5, 6, 8, 10 Gyr. Each panel is 6 Mpc on a side. The cyan dot represents the SMBH of the primary cluster, while the red dot represents the SMBH of the secondary cluster. The primary CC is barely disrupted during merger.
@@
III.3 Heating and Cooling
In this section, we investigate the impact of AGN feedback and radiative cooling on CC destruction in cluster merger simulations. We analyze the evolution of jet power and the accumulated energy injected by AGN feedback for each mcfmcf simulation. These parameters are then compared with the cooling rate and radiative energy to evaluate the balance between heating and cooling mechanisms.
Figure 7: Heating and cooling rate evolution for each mcfmcf simulation. Left: $R=1$ simulations. Center: $R=3$ simulations. Right: $R=10$ simulations. The solid lines are the jet power, representing the heating rate, and the dashed lines are the X-ray luminosity within 50 kpc, representing the cooling rate. The gray lines are from the single cluster simulation for comparison.
Figure 8: Evolution of accumulated heating and cooling energies for each mcfmcf simulation. Left: $R=1$ simulations. Center: $R=3$ simulations. Right: $R=10$ simulations. The solid lines are the accumulated injected AGN energy, and the dashed lines are the accumulated radiative energy. The gray lines are from the single cluster simulation for comparison.
Figure 7 illustrates the evolution of heating and cooling rates, with three subplots, each displaying three cases with the same mass ratio $R$. Simulations with the same impact parameter $b$ are plotted in the same color for comparison. Additionally, we include the result of the single cluster simulation scfscf for reference. In each case, the solid line represents the jet power (i.e., the heating rate), and the dashed line represents the X-ray luminosity within 50 kpc, $L_{x}(<50~\text{kpc})$, (i.e., the cooling rate). The X-ray luminosity $L_{x}(<r)$ is calculated by
$L_{x}(<r)=\int_{0}^{r}n_{e}n_{i}\Lambda(T)4\pi r^{\prime 2}dr^{\prime},$ (12)
@@ III.3 Heating and Cooling
Figure 8 shows the evolution of accumulated injected (i.e., heating) and radiative (i.e., cooling) energies. The injected energy is calculated by multiplying the jet power by each simulation time-step and summing the results. Similarly, the radiative energy is calculated by multiplying the X-ray luminosity within 50 kpc, $L_{x}(<50~\text{kpc})$, by each simulation time-step and summing these energies. The radius of 50 kpc for the cooling rate and radiative energy is chosen to roughly balance the levels of heating rate and injected energy.
@@
For the scfscf simulation, the jet power increases and peaks around 1.8 Gyr to balance with the cooling rate. At this point, the accumulated injected energy approximately balances the radiative energy as well. Subsequently, the heating and cooling rates periodically reach peaks together, indicating that the system is self-regulated. In the merger simulations, the first jet power peak (determined by the initial cooling time) happens to coincide with the first core passage (determined by the initial separation and relative velocity) at $\approx 1.6-2.0$ Gyr. We examine the effect of this timing coincidence by carrying out an additional run of $mcf1$ with an initial separation between the two clusters increased by a factor of 1.5. This approach effectively delays the collision, allowing AGN feedback more time to act. We maintain a constant initial total energy, ensuring that the merging velocity between the two clusters remains consistent with the original $mcf1$ run. In this test, the first core passage occurs at 4.0 Gyr, and the SMBHs merge at 6.2 Gyr, thus delaying the evolution by 2.4 Gyr with respect to $mcf1$. During 2-4 Gyr, the jet power grows strong to balance cooling but does not significantly elevate the central entropy. The evolution of $S_{40}$ is similar to that of $mcf1$, reaching a peak after the first core passage, oscillating during the merger, and gradually decreasing post-merger. At later stages, the NCC is transitioning back to a CC. Therefore, we conclude that the results of $mcf1$ are robust and are not affected by the coincidence between cooling and the first core passage.
@@
For the $mcf1$ simulation, AGN feedback remains quiet after the first core passage until $\approx$ 7 Gyr. During this period, the radiative energy is an order of magnitude higher than the injected energy. This weak AGN feedback suggests that heating from the merger is the dominant mechanism in balancing cooling. As inferred from Figure 2 and Figure 3, the merged cluster reverts to a CC state at later times, and gas mixing induced by the merger is insufficient to further heat the cluster. A significant reason why mixing is inefficient in this particular simulation is its high symmetry; the equal-mass, head-on configuration has the result that the gas components of the two clusters have little opportunity to penetrate each other, resulting in a compressed “pancake” of hot gas in between the two oscillating DM cores (cf. ZuHone, 2011). Only after this feature begins to gravitationally collapse can central densities become large enough for significant cooling. Consequently, the jet power rises again at the later stages to regulate cooling, and the evolution of heating and cooling becomes similar to that of the single cluster case. To ensure the robustness of this result, we conducted an additional simulation introducing a small offset (1% of their relative distance) in the impact parameter. The evolution remains consistent, and the cluster returns to a CC state eventually due to insufficient heating.
@@
For the $mcf5$ simulation, the jet power also remains low after the initial peak at the first core passage. However, as discussed in Section III.2.2, merger heating in this case is insufficient to transform the CC into a NCC, and the AGN is needed to maintain the CC state throughout the simulation. This is supported by Figure 8, where it can be seen that the injected energy from the AGN contributes to a significant fraction to the radiative energy.
For the $mcf2$, $mcf3$, and $mcf6$ simulations, the jet power peaks consistently during 2-4 Gyr. This occurs because in these cases, the disruption from the secondary cluster is not prominent at the first core passage. Consequently, AGN feedback intensifies to balance cooling during this stage. As mentioned in Section III.2.2, the circular motion draws gas into the central region of the primary cluster, maintaining high central density and therefore strong jet power, leading to preheating of the core. In the $mcf3$ and $mcf6$ simulations, the injected energy even exceeds the radiative energy, leading to overheating of the core. After SMBH mergers, the clusters transform into NCCs, and both the heating and cooling rates become low.
For the $R=10$ simulations, the heating and cooling evolution is similar to that of the single cluster case. In the head-on merger ($mcf7$), AGN feedback becomes inactive after the initial strong peak at 1.8 Gyr, as the disruption of the secondary cluster heats the primary core. However, the cluster behaves like a single CC cluster after $\approx$ 8 Gyr. In the off-axis cases, the primary core is barely disrupted by the mergers. The jet power and $L_{x}(<50~\text{kpc})$ quasi-periodically reach peak values, and the injected energy and radiative energy remain balanced at similar levels. This indicates that the cluster is self-regulated by AGN feedback, and the impact of mergers is weak.
IV Discussion
IV.1 Comparison with Previous Works
While there have been no other detailed cluster merger simulations including radiative cooling and AGN feedback, here we compare the results of our mcfmcf simulations with recent simulation results of Valdarnini & Sarazin (2021), which include gas cooling, star formation, and energy feedback from supernovae. Their adiabatic runs are identical to the merger simulations in ZuHone (2011), despite employing the smoothed particle hydrodynamics (SPH) scheme instead of the mesh-based scheme used in the latter. The results of their radiative simulations are generally similar to our mcfmcf simulations in terms of CC destruction. In $R=1$ off-axis cases, the final remnants are NCC clusters. This is due to the continuous gas sloshing and mixing during the later stages of mergers. For the $R=10$ cases, the entropy increase is smaller compared to the merger-only runs, thus the initial CCs are preserved, especially in the off-axis cases.
@@ IV.1 Comparison with Previous Works
IV.2 CC/NCC Transitions
@@
(1) CCs are preserved in minor mergers or mergers that do not trigger sufficient heating. In our $R=10$ simulations, the secondary cluster does not significantly disrupt the primary core, thereby failing to increase central entropy and destroy the CCs. In cases with insufficient heating, such as the $mcf5$ case, heating provided by the merger is insufficient to counteract cooling in the core of the primary cluster. Therefore, AGN feedback is required throughout the simulation, and the injected AGN energy closely balances the radiative energy (Figure 8). Another example of insuffient heating is the $mcf1$ simulation. Although it is heated to a NCC at the time of SMBH merger, the central cooling time is $\sim 2$ Gyr, so that the cluster eventually returns to the CC state at the end of the simulation. In this scenario, the injected AGN energy roughly balances the radiative energy (Figure 8). Therefore, the influence of mergers is subdominant, and AGN feedback plays a critical role for balancing cooling and preventing the clusters from undergoing cooling catastrophe. This is also consistent with the observational evidence that the bulk of CCs appear to be preserved over the past $\sim$ 10 Gyr (McDonald et al. 2017).
@@
(3) CCs are destroyed by the combined effects of mergers and AGN feedback. In major mergers with a large impact parameter (e.g., $mcf2$, $mcf3$, $mcf6$), because direct core collisions occur at later stages, AGN feedback is required to prevent the core from catastrophic cooling. Driven by the mergers, the gas flowing into the primary core further fuels the central SMBH, leading to preheating of the CC. Mixing induced by the infall of the secondary cluster further transforms the core into a NCC state. The injected energy in these cases are comparable or can sometimes exceed the radiative energy. This scenario represents an interesting case where mergers and AGN feedback work hand-in-hand to balance cooling and destroy the CCs.
@@ IV.3 Limitations of the Simulations
V Conclusions
In this work, we conduct binary cluster merger simulations incorporating AGN feedback and radiative cooling to investigate the destruction of CCs. Our primary objective is to assess the impact of cluster mergers and AGN feedback on heating the cluster cores. We implement a purely kinetic jet feedback model mainly coupled with chaotic cold accretion, which has been shown to solve the cooling flow problem (Gaspari et al. 2013), together with an exact cooling scheme into the GAMER code. We conduct three sets of simulations: the m[19]m[1-9] simulations are mergers only; the mc[19]mc[1-9] simulations include mergers with cooling but no AGN feedback; the mcf[19]mcf[1-9] simulations involve mergers with both cooling and AGN feedback. Additionally, we vary the mass ratio and impact parameter to examine the entropy evolution of different merger scenarios. The important findings of our study are summarized below.